OpenStudy (anonymous):

solve using substitution pleasee help answer only is fine x=2y-8 -3x+7y=16

OpenStudy (anonymous):

x=2y-8...Equation 1 -3x+7y=16...equation 2 subtitute eq 1 to eq 2. -3(2y-8)+17y=16 -6y+24+17y=16 11y=16-24 y=-8/11 x=2(-8/11)-8 x=-104/11